PQR Co. has semiannual-pay 6.38 percent coupon bonds with a YTM of 5.51 percent. The current yield on these bonds is 5.96 percent. How many years do these bonds have left until they mature? Can you show step to step how to solve this question? Thanks in advanceQuestion 5 1 pts PQR Co. has semiannual-pay 6.38 percent coupon bonds with a YTM of 5.51 percent. The current yield on these
Answer #1

Bond Price = 63.80/0.0596

Bond Price = $1,070.47

Calculating Time Period,

Using TVM Calculation,

N = [PV = -1,070.47, FV = 1,000, PMT = 31.90, I = 0.0551/2]

N = 21.75

Time Period = 21.75/2 = 10.88 years